OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 8 * 63 = 8 * ? + 8 * 3 = ?

OpenStudy (anonymous):

y are there 2 equal signs?

OpenStudy (anonymous):

NEED HELP PLZ MEDAL AND FAN

OpenStudy (anonymous):

where did you get this question?

OpenStudy (anonymous):

8 * 63 = 8 * ? + 8 * 3 = ? 8 * 63 = ? and from there you can solve for the question mark, but then the middle section just seems unneccassary

OpenStudy (anonymous):

you have to use DISTRIBUTIVE PROPERTY

OpenStudy (anonymous):

to solve the problem

OpenStudy (anonymous):

@acxbox22

OpenStudy (lizz123):

\[8\times63=8\times(60+3)=8\times60+8\times3\]

OpenStudy (anonymous):

=?

OpenStudy (lizz123):

8*20=160 8*3=24 160+24=184